Quatro Co. issues bonds dated January 1, 2017, with a par value of $870,000. The bonds, annual contract rate is 9%, and interest is paid semiannually on June 30 and December 31. The bonds mature in three years. The annual market rate at the date of issuance is 8%, and the bonds are sold for $892,789 1. What is the amount of the premium on these bonds at issuance? 2. How much total bond interest expense will be recognized over the life of these bonds? 3. Prepare an amortization table for these bonds, use the straight-line method to amortize the premium.Semiannual InterestUnamort ized Carrying Value Period-End 01/01/2017 06/30/2017 12/31/2017 06/30/2018 12/31/2018 06/30/2019 12/31/2019 Premium $22,789$ 892,789
Answer #1

1) Premium on bonds payable = 892789-870000 = $22789

2) Interest expense = Interest paid-premium amortization

= (870000*9%*3)-22789

Total interest expense = $212111

3) Amortization

Semi annual interest period Uamortized premium Carrying value
01/01/17 22789 892789
6/30/2017 18891 888991
12/31/2017 15093 885093
06/30/2018 11295 881295
12/31/2018 7497 877497
06/30/2019 3699 873699
12/31/2019 0 870000